# Tiling by Near Coincidence

### v1

@@ IV Dodecagonal Tilings by Near Coincidence
The near-coincidence method provides a versatile framework for varying the size, shape, and orientation of the coincidence window, thereby producing a wide variety of tilings. By exploiting its correspondence with the cut-and-project method, one can reproduce familiar tilings using polygonal coincidence windows that mirror the polygonal acceptance domains of the projection approach. Indeed, dodecagonal tilings generated from a bilayer of two triangular lattices twisted by $30^{\circ}$—shown earlier in Fig. 1b—are related to their cut-and-project counterparts exactly as in the octagonal case. The star-map of Eq. (6) is defined exactly as in Eq. (5), except that now $N=12$, so the ${\mathbf{a}}_{i}$ vectors of Eq. (2) span a pair of triangular rather than square lattices, as illustrated in Fig. 6b (see, for example, Fig. 2 of Oxborrow and Henley [Oxborrow93]).
Figure 8: The three dodecagonal Niizeki–Gähler tilings [Niizeki87, Gahler88], reproduced by the near coincidence of a $30^{\circ}$ twisted triangular bilayer, as presented in Fig. 1b. (a) Three polygonal coincidence windows with dodecagonal symmetry, used for constructing the displayed tilings. (b) The shield tiling, generated with the small blue dodecagon, containing three prototiles: the square, the triangle, and the shield. (c) The Niizeki–Gähler tiling, generated with the nonconvex turquoise 12-fold star, where the shield prototile is replaced by a $30^{\circ}$ rhomb. (d) The Stampfli tiling [Stampfli86], generated with the large dark-gray dodecagon, retaining the same three prototiles as the Niizeki–Gähler tiling.
With this prescription we reproduce the three Niizeki–Gähler dodecagonal tilings [Niizeki87, Gahler88] by applying the near-coincidence method to a $30^{\circ}$ twisted triangular bilayer. The three distinct polygonal coincidence windows used for the construction—equivalent to the polygonal acceptance domains used by Niizeki and Mitani [Niizeki87] and by Gähler [Gahler88]—are shown one on top of the other in Fig. 8a. The smallest window, admitting the lowest density of vertices, is the small blue dodecagon. A slightly larger window is obtained by extending the blue dodecagon into the nonconvex turquoise 12-fold star, formed by the union of a hexagon and its image under a $30^{\circ}$ rotation. The largest window, admitting the highest density of vertices, is obtained by taking the convex closure of the previous star into the large dark-gray dodecagon. The radius of the circle inscribing this large dodecagon is scaled relative to that of the small dodecagon by the factor
$\rho=\frac{2}{\sqrt{3}}\cos{\left(\frac{\pi}{12}\right)}=\sqrt{\frac{2+\sqrt{3}}{3}}\simeq 1.115.$ (8)
See Niizeki and Mitani [Niizeki87] or Gähler [Gahler88] for details.
The resulting tilings are displayed in Figs. 8b8d. As outlined in Fig. 9a, in going from the so-called shield tiling of Fig. 8b to the slightly denser tiling of Fig. 8c, every shield is divided either into a square, a $30^{\circ}$-rhomb, and two triangles (by adding one vertex), or into two triangles and three rhombs (by adding two vertices). Continuing on from this tiling—identified here for convenience as the Niizeki–Gähler tiling—to the densest of the three, shown in Fig. 8d, some of the square–triangle configurations are redivided into a pair of rhombs and a single triangle by adding one vertex. We identify this third tiling as the Stampfli tiling, as it was first described by Stampfli [Stampfli86].
@@
As a final observation, note that the tiling of Fig. 10c, produced using the large dark-gray isotropic coincidence window, still contains many square–triangle pairs that can accommodate an additional vertex. Similarly, all the original shields contain only one or two additional vertices, while there is clearly a third position in each that can accommodate another without violating the single-edge-length preference. Indeed, there exists a range of coincidence thresholds $r$ that increases the density just enough to add all these vertices—but not more—thus producing a tiling without squares, in which all shields are replaced by three rhombs surrounding a 3-fold nonconvex star, which we refer to as the three-arm star, or tristar tile (see Fig. 9c). The resulting tiling, generated with the large yellow coincidence window of Fig. 10a, is shown in Fig. 10d. We note that the same three-arm star tile was observed independently by Sadoc and Impéror-Clerc [Sadoc23] in a different dodecagonal tiling that still contains square tiles.
@@ V Fibonacci tilings by near coincidence
In the previous sections we focused on twisted bilayers with identical length scales, generating both well-known tilings and new variants by means of the near-coincidence method. We now turn to the role of the scaling factor, without introducing any relative rotation, and use the Fibonacci tiling family as our case study. We consider two two-dimensional Fibonacci tilings—one with 4-fold symmetry and the other with 6-fold symmetry—which can be viewed as natural higher-dimensional extensions of the celebrated one-dimensional Fibonacci tiling [Baake24, Jagannathan21]. Both examples are derived from a scaled bilayer construction, in which one layer is rescaled relative to the other by the golden mean $\tau=(1+\sqrt{5})/2$.
V.1 Square Fibonacci Tiling
Figure 11: (a) Two square layers, with the blue scaled by a factor of $\tau$ relative to the red. (b) Two self-similar square Fibonacci tilings [Lifshitz02], obtained with a $\tau$-scaled pair of centered square coincidence windows, revealing four different orientations of the canonical substitution rules [Lifshitz02, Fig. 3]. (c) Two self-similar square Fibonacci tilings, obtained with a $\tau$-scaled pair of shifted square coincidence windows, revealing the canonical Fibonacci substitution rules, $S\rightarrow L$ and $L\rightarrow LS$, in both horizontal and vertical directions.
We begin with the square Fibonacci tiling [Lifshitz02], obtained using a $\tau$-scaled square bilayer, shown in Fig. 11a. Empirical inspection identifies the appropriate choice for the coincidence window as being square, producing the tiling displayed with thick blue lines in Fig. 11b. As expected in a $\tau$-scaled bilayer, one needs a minimum of two edge lengths in order properly to connect the vertices. Indeed, visual inspection of the tiling reveals the familiar Fibonacci sequence of long ($L$) and short ($S$) edges in both vertical and horizontal directions. Accordingly, the tiling consists of three prototiles: a small $S\times S$ square, a large $L\times L$ square, and an $S\times L$ rectangle, appearing in two orientations. Thus, the near-coincidence method seems to work well in producing the intended tiling, yet as we now show, it does not map directly neither to the standard cut-and-project procedure nor to the canonical substitution rules, used to generate the tiling, as we have seen so far in all previous examples.
Figure 12: The standard star maps used in the cut-and-project method for the (a) square; and the (b) hexagonal Fibonacci tilings.
Indeed, the mapping of the near-coincidence construction to that of the cut-and-project method is not as straightforward as in the previous examples with twisted bilayers. The reason for this is that the cut-and-project star map, given by the basis vectors in Fig. 12a, not only flips the direction of the blue vectors, but also interchanges the lengths of the red and blue vectors, so that
$\mathbf{b}_{0}=\mathbf{a}_{1},\quad\mathbf{b}_{1}=-\mathbf{a}_{0},\quad\mathbf{b}_{2}=\mathbf{a}_{3},\quad\mathbf{b}_{3}=-\mathbf{a}_{2},$ (9)
where here we have ${\mathbf{a}}_{1}=\tau{\mathbf{a}}_{0}$ and ${\mathbf{a}}_{3}=\tau{\mathbf{a}}_{2}$. Compare with Eq. (5).
To see the consequence of this behavior, let us concentrate on the one-dimensional horizontal axis for ease of exposition. Following Eq. (9), the star-map condition in one dimension takes the form
$\sum_{k=0}^{1}n_{k}{\mathbf{b}}_{k}=n_{0}{\mathbf{a}}_{1}-n_{1}{\mathbf{a}}_{0}=\left(n_{0}\tau-n_{1}\right){\mathbf{a}}_{0}\in\text{\ AD}.$ (10)
It is known to produce the Fibonacci tiling via the cut-and-project method, with a canonical acceptance domain whose length $|\text{AD}|=1+\tau=\tau^{2}$, accepting a tiling vertex at the position $\left(n_{0}+n_{1}\tau\right){\mathbf{a}}_{0}$, for every $(n_{0},n_{1})$ pair satisfying the condition in Eq. (10). In contrast, using the same condition, but viewing the acceptance domain AD as a coincidence window CW of the same length, and applying the near-coincidence method, selects the same $(n_{0},n_{1})$ pairs, but places tiling vertices at the midpoint positions $\left(n_{0}\tau+n_{1}\right){\mathbf{a}}_{0}/2$ instead. One thus obtains a correct Fibonacci sequence of the two edge lengths, albeit with a “short” edge of length $\tau/2$, and a “long” edge of length $1/2$.
One can remedy this situation by successively shrinking the dimension of the coincidence window by factors of $\tau$, thereby affecting substitution rules and obtaining sparser Fibonacci tilings. The fact that $\tau$ is a Pisot number—an eigenvalue greater than unity of a substitution matrix whose remaining eigenvalue $1/\tau$ is less than unity in magnitude—ensures that successive applications of the substitution to almost any initial vector of edge lengths, approach the eigenvector corresponding to the eigenvalue $\tau$. The edge-length ratio thus tends to the canonical ratio of $\tau\!\!:\!\!1$ between the sparser long and short edges, even though the initial “long” and “short” edges have the opposite length ratio of $1\!\!:\!\!\tau$. Figure 13 demonstrates this behavior in the one-dimensional case for a sequence of nine substitutions. In this figure, a centered coincidence window is used, defined by the condition
$|n_{0}\tau-n_{1}|<\frac{1}{2}(1+\tau).$ (11)
Thus, a practical rendering of the two-dimensional square tiling, using near coincidence of scaled square bilayers, can achieve any required accuracy of the canonical edge-length ratio $\tau$ by taking sufficiently small coincidence windows. Alternatively, one can simply follow the sequence of “long” and “short” edges, obtained with the canonical coincidence window, but render the tiling with correctly scaled edges. This observation holds as well for the hexagonal Fibonacci tiling, discussed below.
@@ V.1 Square Fibonacci Tiling
Figure 13: Successive one-dimensional Fibonacci tilings for centered coincidence windows of diminishing size $\tau^{2-n}$, shown row by row and labeled along the vertical axis. Each row displays the positions of the tiling vertices admitted by the $n^{th}$ coincidence window, with curves connecting points of the same index across successive orders. For all $n$, the sequence $l_{n}s_{n}l_{n}s_{n}l_{n}l_{n}s_{n}l_{n}$… exhibits the correct Fibonacci order, yet with a noncanonical edge-length ratio $l_{n}/s_{n}$ that eventually approaches $\tau$ with increasing order, as detailed in the upper-left inset. The upper-right inset demonstrates that at order $n=0$, with the largest coincidence window, one even has $l_{0}<s_{0}$, where $l_{0}$ labels the short edges and $s_{0}$ labels the long ones.
A direct attempt to recover the canonical Fibonacci substitution rules: $S\rightarrow L$ and $L\rightarrow LS$—or its two-dimensional generalization [Lifshitz02, Fig. 3]—by overlaying a pair of $\tau$-scaled Fibonacci tilings, fails as well. Instead, one obtains alternative symmetric substitution rules  [^1] Alternative, or conjugate, substitution rules that define the same tiling are discussed elsewhere. See, for example, the discussion of metallic-mean sequences in Baake and Grimm [Baake13, Sec. 4.4]. for the $L$ tiles, whereby every $LL$ pair is substituted by $SLLS$, and individual $L$ tiles are replaced by either $LS$ or $SL$ with equal probability, depending on their surroundings. This can be seen by comparing any pair of successive one-dimensional tilings in Fig. 13, or the pair of $\tau$-scaled square tilings in Fig. 11b. The use of such substitution rules, while correct, is less convenient, as one needs to keep track of the orientations of the $L$ tiles in each iteration of the substitution. Compare with the use of red arrows in the Ammann-Beenker tiling above (Fig. 7). An encoding of the orientation can be achieved by distinguishing two types of $L$ and $S$ tiles using overhead arrows and removing the arrows at the end, with substitution rules given by $\overset{{\shortrightarrow}}{L}\to\overset{{\shortrightarrow}}{L}\overset{{\shortrightarrow}}{S}$, $\overset{{\shortleftarrow}}{L}\to\overset{{\shortleftarrow}}{S}\overset{{\shortleftarrow}}{L}$, $\overset{{\shortrightarrow}}{S}\to\overset{{\shortleftarrow}}{L}$, and $\overset{{\shortleftarrow}}{S}\to\overset{{\shortrightarrow}}{L}$.
This situation arises because we use a symmetric coincidence window, measuring the difference vector ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{p}_{1}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{p}_{2}}$ with respect to the center of the coincidence interval (11) in one-dimension, or to the center of the square coincidence window in two dimensions. If instead, we measure the difference vector with respect to the left edge of the coincidence interval or bottom-left corner of the square window, we obtain the tiling of Fig. 11c. In doing so, we admit nearly coincident blue-red pairs and reject nearly coincident red-blue pairs, thereby breaking the symmetry of the coincidence window. Such anisotropy may naturally arise in physical systems where the interaction between layers is directionally dependent. As expected, the shifted coincidence window breaks the symmetry of the substitution, globally replacing $L$ by either $LS$ or $SL$, but the choice turns out to depend on the parity of the substitution order $n$. These alternating substitution rules can be described, again, by employing oriented $L$ and $S$ tiles, as follows: $\overset{{\shortrightarrow}}{L}\to\overset{{\shortleftarrow}}{L}\overset{{\shortleftarrow}}{S}$, $\overset{{\shortleftarrow}}{L}\to\overset{{\shortrightarrow}}{S}\overset{{\shortrightarrow}}{L}$, $\overset{{\shortrightarrow}}{S}\to\overset{{\shortleftarrow}}{L}$, and $\overset{{\shortleftarrow}}{S}\to\overset{{\shortrightarrow}}{L}$.
V.2 Hexagonal Fibonacci Tiling
Figure 14: The $H_{00}$ Hexagonal Fibonacci tiling [Coates24]. (a) A $\tau$-scaled honeycomb bilayer. (b) The hexagonal Fibonacci tiling, with excess points marked by empty circles. (c) The tiling after removing the excess points, with vertices color-coded according to their height values, as defined in Eq. (12). (d) The four coincidence windows corresponding to the four vertex types, with the points shown in gray.
@@ V.2 Hexagonal Fibonacci Tiling
We next consider the near coincidence of a $\tau$-scaled honeycomb bilayer, shown in Fig. 14a. In this case the optimal shape of the coincidence window is not obvious. Fortunately, a default circular window leads to the tiling shown in Fig. 14b, which contains excess points—already shown as empty black circles—that are easy to remove by simple local rules. These excess points are either dangling vertices that connect to only one other vertex, or redundant vertices that lie along already existing edges. The clean tiling, with no excess points, shown in Fig. 14c is the $H_{00}$ hexagonal Fibonacci tiling of Coates et al. [Coates24].
@@
Mapping the accepted and rejected vertices into the circular coincidence window, as we did in the octagonal case (Fig. 5), reveals a central hexagon containing only accepted vertices. The central hexagon is surrounded by mixed regions, which in turn are surrounded by an outer perimeter containing only rejected vertices. This suggests the existence of several coincidence windows, for different subsets of vertices. Following the construction of Coates et al. [Coates24], one indeed finds that four coincidence windows are required, as the cut-and-project construction of this tiling is from four different two-dimensional cuts through a six-dimensional embedding space (for details see Ref. [Coates24]).
Given a vertex with coordinates $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}n_{0},n_{2},n_{4}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{1},n_{3},n_{5}})$ with respect to the basis vectors in Fig. 12b, one assigns it a color based on a pair of heights
$({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}h_{s}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h_{l}})=({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}n_{0}+n_{2}+n_{4}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{1}+n_{3}+n_{5}})\bmod 3,$ (12)
@@
which take the values (1,1) [red], (1,2) [orange], (2,1) [green], and (2,2) [blue]. The tiling in Fig. 14c is colored according to this scheme, and so are the corresponding points mapped onto the coincidence window in Fig. 14d. This partitioning of the vertices yields a pair of oppositely oriented orange and green triangular coincidence windows, and a pair of oppositely oriented red and blue 3-fold symmetric hexagonal coincidence windows, thus maintaining the overall hexagonal symmetry.

### v2

@@ IV Dodecagonal Tilings by Near Coincidence
[…]
Figure 8: The three dodecagonal Niizeki–Gähler tilings [5, 6], reproduced by the near coincidence of a $30^{\circ}$ twisted triangular bilayer, as presented in Fig. 1b. (a) Three polygonal coincidence windows with dodecagonal symmetry, used for constructing the displayed tilings. (b) The shield tiling, generated with the small blue dodecagon, containing three prototiles: the square, the triangle, and the shield. (c) The Niizeki–Gähler tiling, generated with the nonconvex turquoise 12-fold star, where the shield prototile is replaced by a $30^{\circ}$ rhomb. (d) The Stampfli tiling [4], generated with the large dark-gray dodecagon, retaining the same three prototiles as the Niizeki–Gähler tiling.
With this prescription we reproduce the three Niizeki–Gähler dodecagonal tilings [5, 6] by applying the near-coincidence method to a $30^{\circ}$ twisted triangular bilayer. The three distinct polygonal coincidence windows used for the construction—equivalent to the polygonal acceptance domains used by Niizeki and Mitani [5] and by Gähler [6]—are shown one on top of the other in Fig. 8a. The smallest window, admitting the lowest density of vertices, is the small blue dodecagon. A slightly larger window is obtained by extending the blue dodecagon into the nonconvex turquoise 12-fold star, formed by the union of a hexagon and its image under a $30^{\circ}$ rotation. The largest window, admitting the highest density of vertices, is obtained by taking the convex closure of the previous star yielding the large dark-gray dodecagon.
The resulting tilings are displayed in Figs. 8b8d. As outlined in Fig. 9a, in going from the so-called shield tiling of Fig. 8b to the slightly denser tiling of Fig. 8c, every shield is divided either into a square, a $30^{\circ}$-rhomb, and two triangles (by adding one vertex), or into two triangles and three rhombs (by adding two vertices). Continuing on from this tiling—identified here for convenience as the Niizeki–Gähler tiling—to the densest of the three, shown in Fig. 8d, some of the square–triangle configurations are redivided into a pair of rhombs and a single triangle by adding one vertex. We identify this third tiling as the Stampfli tiling, as it was first described by Stampfli [4].
@@
As a final observation, note that the tiling of Fig. 10c, produced using the large dark-gray isotropic coincidence window, still contains many square–triangle pairs that can accommodate an additional vertex. Similarly, all the original shields contain only one or two additional vertices, while there is clearly a third position in each that can accommodate another without violating the single-edge-length preference. Indeed, there exists a range of coincidence thresholds $r$ that increases the density just enough to add all these vertices—but not more—thus producing a tiling without squares, in which all shields are replaced by three rhombs surrounding a 3-fold nonconvex star, which we refer to as the three-arm star, or tristar tile (see Fig. 9c). The resulting tiling, generated with the large yellow coincidence window of Fig. 10a, is shown in Fig. 10d. We note that the same three-arm star tile was observed independently by Sadoc and Impéror-Clerc [27] in a different dodecagonal tiling that still contains square tiles.
@@ V Fibonacci tilings by near coincidence
In the previous sections we focused on twisted bilayers with identical length scales, generating both well-known tilings and new variants by means of the near-coincidence method. We now turn to the role of the scaling factor, without introducing any relative rotation, and use the Fibonacci tiling family as our case study. We begin with the one-dimensional Fibonacci tiling [28, 29], which captures the essential features of the construction, and then show how it extends naturally to two-dimensional square [19] and hexagonal [20] Fibonacci tilings. Both examples arise from a scaled bilayer construction, in which one layer is rescaled relative to the other by the golden mean $\tau=(1+\sqrt{5})/2$.
V.1 One-dimensional Fibonacci tilings from scaled bilayers
To expose the essential features of the scaled bilayer construction, we first consider the one-dimensional case. The cut-and-project construction of the one-dimensional Fibonacci tiling may be formulated as a projection of the square lattice $\mathbb{Z}^{2}$, with physical-space coordinates $(n_{0}+n_{1}\tau)$ and internal-space coordinates $(n_{0}\tau-n_{1})$. The acceptance condition selects those integer pairs $(n_{0},n_{1})$ for which
$n_{0}\tau-n_{1}\in\text{AD},$ (8)
with a canonical acceptance domain whose length $|\text{AD}|=1+\tau=\tau^{2}$, yielding the standard Fibonacci tiling with edge lengths in the ratio $\tau\!:\!1$.
Within the near-coincidence framework, the same integer pairs are selected by viewing the acceptance domain AD as a coincidence window CW of the same size. However, the tiling vertices are placed at the midpoint positions $(n_{0}\tau+n_{1})/2$, rather than at $(n_{0}+n_{1}\tau)$. As a result, one obtains the correct Fibonacci ordering of long and short edges, but not their canonical length ratio: the “short” and “long” edges initially appear with lengths $\tau/2$ and $1/2$, respectively.
This discrepancy can be resolved in two ways. One may simply retain the correct sequence of long and short edges and rescale their lengths by hand to enforce the canonical ratio. Alternatively, one may achieve any required accuracy by shrinking the coincidence window by successive factors of $\tau$, thereby generating sparser tilings whose edge-length ratio approaches $\tau$. This occurs because $\tau$ is a Pisot number—an eigenvalue greater than unity of a substitution matrix whose remaining eigenvalue $1/\tau$ is less than unity in magnitude—ensuring that successive substitutions drive the edge-length ratio toward its canonical value, even when starting from the inverted ratio. This behavior is illustrated in Fig. 11, where successive coincidence windows of diminishing size are used.
@@ V.1 One-dimensional Fibonacci tilings from scaled bilayers
Figure 11: Successive one-dimensional Fibonacci tilings for centered coincidence windows of diminishing size $\tau^{2-n}$, shown row by row. The correct Fibonacci order is obtained for all $n$, while the edge-length ratio $l_{n}/s_{n}$ approaches $\tau$ only with increasing order, as detailed in the upper-left inset. The upper-right inset demonstrates that at order $n=0$, with the largest coincidence window, one begins with the inverted ratio relative to the canonical one, where $l_{0}$ labels short edges and $s_{0}$ labels long ones.
[…]
V.2 Square Fibonacci tiling
The two-dimensional square Fibonacci tiling [19] follows immediately as a direct product extension of the one-dimensional construction. Using a $\tau$-scaled square bilayer (Fig. 12a) and a square coincidence window, one obtains the tiling shown in Fig. 12b. The tiling consists of three prototiles: a small $S\times S$ square, a large $L\times L$ square, and an $S\times L$ rectangle, with the Fibonacci sequence appearing independently along both the horizontal and the vertical directions.
Figure 12: Square Fibonacci tilings obtained from a $\tau$-scaled bilayer: (a) the scaled layers; (b) a pair of tilings, obtained using centered $\tau$-scaled coincidence windows, which lead to symmetric substitution rules, revealing all four orientations of the canonical substitution rules [19, Fig. 3]; (c) using shifted coincidence windows yields the canonical Fibonacci substitution with alternating orientation.
The corresponding cut-and-project star-map, shown in Fig. 13a, as well as the distinction between centered and shifted coincidence windows, carry over directly from the one-dimensional case. Here as well, one observes the same discrepancy where the edge-length ratio approaches its canonical value only as the coincidence window shrinks; and centered coincidence windows lead to symmetric substitution rules as shown in Fig. 12b, while shifted windows recover the canonical Fibonacci substitution, as illustrated in Fig. 12c.
Figure 13: The standard star maps used in the cut-and-project method for the (a) square; and (b) hexagonal Fibonacci tilings.
V.3 Hexagonal Fibonacci tiling
@@
We next consider the near coincidence of a $\tau$-scaled honeycomb bilayer, shown in Fig. 14a. In this case the optimal shape of the coincidence window is not obvious. Nevertheless, a natural default choice is to use a sufficiently small circular coincidence window, ensuring (to any desired accuracy) the canonical edge-length ratio, as discussed in the previous section, without affecting the hexagonal symmetry. This produces the tiling in Fig. 14b, containing excess points, marked by empty circles, that are easily removed by simple local rules: eliminating dangling vertices that connect to only one other vertex, and redundant vertices that lie along existing edges. The resulting clean tiling, shown in Fig. 14c, is precisely the $H_{00}$ hexagonal Fibonacci tiling of Coates et al. [20].
@@
Mapping the accepted and rejected vertices onto the circular coincidence window, shown in Fig. 14d, reveals a central hexagon containing only accepted vertices, surrounded by various mixed regions containing both accepted and rejected vertices, which in turn are surrounded by an outer perimeter containing only rejected vertices. Further inspection suggests that four coincidence windows are required, corresponding to four distinct subsets of vertices. These depend on the parities of the numbers of small red steps and large blue steps—indicated in Fig. 13b—required to reach a vertex from the origin. This is in agreement with the cut-and-project construction of this tiling, which requires four different two-dimensional cuts through a six-dimensional embedding space (for details see Ref. [20]).
@@
The tiling in Fig. 14c is colored according to this scheme, and so are the corresponding points mapped onto the coincidence window in Fig. 14d. This partitioning of the vertices yields a pair of oppositely oriented orange and green triangular coincidence windows, and a pair of oppositely oriented red and blue 3-fold symmetric hexagonal coincidence windows, thus maintaining the overall hexagonal symmetry.
Thus, in contrast to the square case, the hexagonal Fibonacci tiling involves a richer structure, with multiple vertex classes and multiple coincidence windows, reflecting the higher-dimensional nature of the corresponding cut-and-project construction.
Figure 14: The $H_{00}$ Hexagonal Fibonacci tiling [20]. (a) A $\tau$-scaled honeycomb bilayer. (b) The hexagonal Fibonacci tiling, with excess points marked by empty circles. (c) The tiling after removing the excess points, with vertices color-coded according to their parity classes, as described in the text. (d) The four coincidence windows corresponding to the four vertex types, with rejected points shown in gray.